OpenStudy (anonymous):

Jason has four lengths of rope. The lengths are 3 inches, 15 centimeters, 0.3 meters, and 0.5 feet. About how much rope does Jason have? [1 inch ≈ 2.5 cm] A) 20 inches B) 23 inches C) 27 inches D) 32 inches

OpenStudy (anonymous):

Convert them into inches and add them all up.

OpenStudy (lukebluefive):

\[15 cm \approx \frac{15cm}{2.5inches} \approx 6 inches\] \[0.3 meters = 0.3 meters \times 100 \frac{cm}{meter} = 30 cm \approx 12 inches\] Can you handle the rest?

OpenStudy (anonymous):

i think so and jason how do i convert them into inches

OpenStudy (anonymous):

Jazzy, 1 inch = 2.5 cms. Now, how much is 15 cms? It is 15/2.5 which is 6inches

OpenStudy (anonymous):

hold up..let me do the math rite quick plzz

OpenStudy (anonymous):

Similarly, 1 meter = 100 cms. Then how many cms is 0.3 meters? It is 0.3*100 which is 30 cms. Now, convert the cms into inches. 1 inch = 2.5 cms. How many inches is 30 cms? 30/2.5 = 12inches.

OpenStudy (anonymous):

ok i think i get it now